### v2

@@ 5 Conclusion
In this work, we propose graph-based data selection methods that build the train set before the model’s training, such that the risk is alleviated. We develop two methods, GraphReach and MaxSpec, that build the gene perturbation train set using the input knowledge graph. Specifically they select the perturbations that are expected to improve the GNN’s generalization. In the experiments, we see that the data-selection methods perform better then active learning when we have a few hundreds of perturbations, while active learning performs better when we have a few tens. That said GraphReach is the most robust among all the methods, achieving first or second performance in both datasets and both metrics. Taking into consideration these results and that data-selection is devoid of the aforementioned risk, we deem GraphReach a valid approach to train effectively a gene-expression model.
In the future we plan to examine hybrid techniques that utilize graph-based data-selection methods to jump-start the training up to a point and continue with active learning runs for the rest. Moreover, we plan to experiment with multi-gene perturbations and potentially use more involved selection criteria that are common in graph active learning [44, 45].
@@ Appendix 0.A The GEARS model
$L=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{K}\sum^{K}_{u=1}(Y_{u}-\hat{Y}_{u})^{(2+\gamma)}$ (16)

### v3

@@ 5 Conclusion
In this work, we propose graph-based data filtering methods that build the train set before the model’s training, such that the risk is alleviated. We develop two methods, GraphReach and MaxSpec, that build the gene perturbation train set using the input knowledge graph. Specifically they select the perturbations that are expected to improve the GNN’s generalization. In the experiments, we see that the data filtering methods perform better then active learning when we have a few hundreds of perturbations, while active learning performs better when we have a few tens. That said GraphReach is the most robust among all the methods, achieving first or second performance in both datasets and both metrics. Taking into consideration these results and that data filtering is devoid of the aforementioned risk, we deem GraphReach a valid approach to train effectively a gene-expression model.
In the future we plan to examine hybrid techniques that utilize graph-based data filtering methods to jump-start the training up to a point and continue with active learning runs for the rest. Moreover, we plan to experiment with multi-gene perturbations and potentially use more involved selection criteria that are common in graph active learning [45, 47].
@@ Appendix 0.A The GEARS model
$L=\frac{1}{T}\sum^{T}_{k=1}\frac{1}{T_{k}}\sum^{T_{k}}_{l=1}\frac{1}{K}\sum^{K}_{u=1}(Y_{u}-\hat{Y}_{u})^{2}$ (16)